Grade 11Inorganic Chemistry

4.4 g of CO2 and 2.24 litre of H2 at STP are in a container. The total number of molecules in the container will be?

Hialo Sharon
8 Years agoGrade 11

2 Answers

Mahima Kanawat
8 Years ago
Dear student... Moles of CO2=4.4/44=0.1, moles of H2=2.24/22.4=0.1 ; therefore total moles of gases =0.2 and molecules are 0.2 × Na = 0.2×6×10²³=1.2×10²³....WITH REGARDS, MAHIMA (ASKIITIANS FORUM EXPERT)
Unni
7 Years ago
Always remember no:of molecules=n×Na where n=no:of moles =(1)
And at STP ,n=volume of gas in l /22.4
After finding total moles n=0.2 put in =(1)
No of moles of co2 =given mass/molar mass
                                 =4.4/44=0.1
No of moles ofH2=2.24/22.4=0.1
Total=0.1+0.1=0.2